(Factorise) Find the difference of two squares:
16x^2-25y^2
Please show all working

Found 2 solutions by Edwin McCravy, Theo:
Answer by Edwin McCravy(20081)

First learn the principle of factorising 
the difference of two squares. Here is 
why it works:

A²-B² = (A-B)(A+B)

Proof: (A-B)(A+B) = A²+AB-AB-B² = A²-B²

So we memorise:

A²-B² is factorized as (A-B)(A+B)

--------------------------------------

16x²-25y²

Since 16=4² and 25=5², we can rewrite that as

4²x²-5²y²

and again as

(4x)²-(5y)²

Since we have memorized that

A²-B² is factorised as (A-B)(A+B)

We can substitute 4x for A and 5y for B and get that

(4x)²-(5y)² is factorised as (4x-5y)(4x+5y)

Edwin


Answer by Theo(13342)

16x^2 - 25y^2 is equal to:

(4x)^2 - (5y^2) which is equal to:

(4x-5y) * (4x + 5y)